Solve the following problems on Income elasticity of demand and classify the nature of elasticity and good. 3)Suppose the initial income of a consumer is Rs.2000/- per week and he purchases 20 units of a commodity. In the next month the income of the consumer rises to Rs.3000/- per week, as a result the demand rises to 40 units. 4) Suppose the initial income of a consumer is Rs.2000/- per week and he purchases 20 units of a commodity. In the next month the income of the consumer rises to Rs.3000/- per week, as a result the demand rises to 25 units. 5) Suppose the initial income of a consumer is Rs.2000/- per week and he purchases 20 units of a commodity. In the next month the income of the consumer rises to Rs.3000/- per week, as a result the demand falls to 10 units.
1
Expert's answer
2021-10-26T14:45:17-0400

Income elasticity of demand (YED)="\frac{\%change in Q}{\%change in P}"


(3)


(YED)="\frac{\%change in Q}{\%change in P}"="\frac{\frac{40-20}{40}\times 100\%}{\frac{1000}{3000}\times100\%}=1.5"


Since the value is positive and greater than 1, the good is highly elastic. This good is a luxury good since its demand increases as soon as it becomes affordable.



(4)


(YED)="\frac{\%change in Q}{\%change in P}"="\frac{\frac{25-20}{25}\times 100\%}{\frac{1000}{3000}\times100\%}=0.6"


A change in the income has not changed the demand of the good substantially therefore the good is inelastic.This good is a basic good since its demand is less sensitive to a change in income


(5)


(YED)="\frac{\%change in Q}{\%change in P}"="\frac{\frac{10-20}{20}\times 100\%}{\frac{1000}{3000}\times100\%}=-1.5"



A change in the income has changed the demand of the good negatively and therefore the good is inelastic.This good is an inferior good since its demand a rise in income has led to a decrease in demand of the good.